# Quantum teleportation implies symmetry-protected topological order

### v2

@@ 4.1 Action of dilated channels
Here we consider the action of single-qubit measurements (2.3) and “Clifford” recovery operations (34) on the logical operators (2.2) in canonical-form protocols. We first derive an effective recovery channel for the logical operators $X_{{\rm L}}$ and $Z_{{\rm L}}$ in Lemma 7. We then consider the combination of measurements and feedback on the logical operators in Lemma 8. Both proofs extend to any $k\geq 1$.
We now present Lemma 7, which establishes an effective recovery channel $\mathcal{R}_{\rm eff}$ for any $\mathcal{R}$ in a canonical-form standard teleportation protocol (see Defs. 4 and 6).
@@
$\mathcal{W}\cong\mathcal{W}_{\rm eff}=\mathcal{R}_{x}\,\mathcal{R}_{z}\,\mathcal{M}\,\mathcal{U}\,,~{}~{}$ (37)
where $\mathcal{R}_{x}$ ($\mathcal{R}_{z}$) act nontrivially only on $X_{{\rm L}}$ ($Z_{{\rm L}}$) as
$\mathcal{R}_{x}$ $=\mathds{1}\otimes\widetilde{P}^{(0)}_{r_{x}}+Z_{{\rm L}}\otimes\widetilde{P}^{(1)}_{r_{x}}$ (38a)
$\mathcal{R}_{z}$ $=\mathds{1}\otimes\widetilde{P}^{(0)}_{r_{z}}+X_{{\rm L}}\otimes\widetilde{P}^{(1)}_{r_{z}}\,,~{}~{}$ (38b)
where $r_{\nu}\subset\widetilde{V}$ is a subset of Stinespring qubits determined by $\mathcal{R}$ (34), the projectors correspond to outcome parities of the cluster $r_{\nu}$ (35), $\mathcal{R}_{x}$ and $\mathcal{R}_{z}$ effectively commute, and $\mathcal{R}_{\nu}=\mathds{1}$ if $\mathcal{R}$ acts trivially on $\sigma^{\nu}_{\rm L}$.
The proof of Lemma 7 appears in App. E. The Lemma establishes (i) when and how a Clifford recovery operation $\mathcal{R}_{s}$ (34) acts nontrivially on a logical operator $\Gamma$, (ii) that the effective recovery channel $\mathcal{R}_{\rm eff}=\mathcal{R}_{x}\,\mathcal{R}_{z}$ (37) is equivalent to $\mathcal{R}$ in its action on all logicals $\Gamma$ (but not other operators), and (iii) that the recovery channels for different logical operators can be written in any order, and do not mix. For $k>1$ logical qubits, the pair of channels $\mathcal{R}_{\nu}$ (37) are replaced by $2k$ channels $\mathcal{R}_{n,\nu}$ for $n\in[1,k]$, in any order. We stress that both $\mathcal{M}$ and $\mathcal{U}$ in $\mathcal{W}$ (37) are unaltered compared to $\mathcal{W}=\mathcal{R}\,\mathcal{M}\,\mathcal{U}$ (31).
@@
Consider a standard teleportation protocol $\mathcal{W}$ in canonical form (see Defs. 4 and 6) involving the measurement of $M$ (single-qubit) observables $A_{j}$ (32). Suppose that $\Gamma$ is a logical operator (e.g., $X_{{\rm L}}$) acting on the final state $\lvert\Psi_{T}\rangle$ (7). Then $\mathcal{W}$ fails to achieve teleportation (1) unless
$\left[A_{j},\,\Gamma(t_{j})\right]=0\,,~{}~{}$ (39)
@@
for all observables $A_{j}$ (32) and all logical operators $\Gamma$ evolved to the Heisenberg time $t_{j}$ immediately prior to the measurement of $A_{j}$. Moreover, if $\mathcal{R}$ acts nontrivially on $\Gamma$, then the measurement channel acts as
$\Gamma(t_{j-1})=\mathcal{M}^{\dagger}_{j}\,\Gamma(t_{j})\,\mathcal{M}_{j}\cong\bar{A}_{j}\,\Gamma(t_{j})\,,~{}~{}$ (40)
@@
and trivially otherwise, where $\bar{A}_{j}$ is the involutory part (4) of $A_{j}$ (32) and $\cong$ becomes a genuine equality upon projecting onto the default initial Stinespring state $\lvert 0\rangle$.
@@
The proof of Lemma 8 appears in App. F. Simply put, Lemma 8 extends Lemma 7 to the combination of a measurements and recovery operations acting on involutory logical operators $\Gamma\in\boldsymbol{\Gamma}\subset\mathbf{P}_{*}{\rm ph}$. For a given $\Gamma$, this combination either acts trivially, violates the assumption that $\Gamma(t)$ remains a logical operator (i.e., involutory), or attaches the involutory part $\bar{A}_{j}$ (4) of the measured operator $A_{j}$ to $\Gamma(t_{j})$, where $t_{j}$ is the (Heisenberg) time immediately prior to the $j$th measurement. Moreover, $A_{j}$ must commute with $\Gamma(t_{j})$. For $\mathcal{M}_{j}$ to act nontrivially on $\Gamma$, an odd number of recovery Paulis $\chi_{j,k}$ conditioned on the outcome $j$ must anticommute with the operators $\Gamma(\tau_{j,k})$, where $\tau_{j,k}$ is the (Heisenberg) time immediately prior to the channel $\mathcal{R}_{j,k}$. This holds for each measured observable $A_{j}$ and all logical operators $\Gamma$.
@@
Note that Lemma 8 holds for arbitrary (i.e., non-Clifford) $\mathcal{U}$ in the canonical-form protocol (31). The proof of Lemma 8 holds independently of any physical operations that act prior to the first measurement. Additionally, any physical unitary can be “pulled through” a measurement $\mathcal{M}_{j}$ without affecting the result of Prop. 4, so that Lemma 8 continues to hold. Likewise, one can safely “pull through” any recovery operations $\mathcal{R}$ that act prior to some measurement (in the Schrödinger picture), to realize a canonical-form protocol (31).
@@
Summarizing the above, the action of any canonical-form standard teleportation protocol $\mathcal{W}$ on a logical operator $\Gamma\in\boldsymbol{\Gamma}$ is constrained by Lemmas 7 and 8. The former establishes that $\mathcal{R}$ is equivalent in its action on all $\Gamma\in\boldsymbol{\Gamma}$ to the same protocol $\mathcal{W}$ with $\mathcal{R}\to\mathcal{R}_{\rm eff}=\prod_{n=1}^{k}\,\mathcal{R}_{n,x}\,\mathcal{R}_{n,z}$ (37), where $\chi_{n,\nu}$ is the other logical operator for qubit $n$ or the identity. The latter establishes that each measured observable $A_{j}$ must commute with every logical operator $\Gamma(t_{j})$ at the Heisenberg time $t_{j}$ immediately prior to $\mathcal{M}_{j}$, and that $\mathcal{M}_{j}$ attaches $\bar{A}_{j}$ (4) to $\Gamma(t_{j})$ if the effective channel $\mathcal{R}_{\Gamma}$ (7) is nontrivial (i.e., $\chi_{\Gamma}\neq\mathds{1}$).
@@ 4.2 At least two measurements are required
Consider a physical teleportation protocol $\mathcal{W}=\mathcal{R}\mathcal{M}\mathcal{U}$ (31) in canonical form (see Defs. 3 and 4), where $\mathcal{U}$ captures generic time evolution on $\mathcal{H}_{\rm ph}$ (1) for duration $T$ and obeys the bound (28) with Lieb-Robinson velocity $v$; $\mathcal{M}$ realizes exactly one projective measurement of some (effectively) involutory observable $A$ (4); and $\mathcal{R}$ is an arbitrary recovery channel conditioned on the measurement’s outcome. If the protocol $\mathcal{W}$ teleports a single logical qubit a distance $L$ in time $T$, then $L\leq vT$.
The proof of Prop. 9 appears in App. G. Prop. 9 establishes that a single projective measurement cannot lead to a speedup compared to a protocol $\mathcal{W}^{\prime}$ with no measurements at all. Essentially, Prop. 1 is only fulfilled with $L>vT$ if the dilated channels attach distinct operators (via Lemma 8) to, e.g., the logical operators $X_{{\rm L}}$ and $Z_{{\rm L}}$; otherwise, $Y_{{\rm L}}$ can only travel a distance $L\leq vT$.
@@
Consider a standard teleportation protocol $\mathcal{W}$ in canonical form $\mathcal{W}=\mathcal{R}\,\mathcal{M}\,\mathcal{U}$ (see Defs. 4 and 6), where $\mathcal{U}$ realizes unitary time evolution on $\mathcal{H}_{\rm ph}$ for duration $T$ and obeys the bound (28), $\mathcal{M}$ involves two projective measurements, and $\mathcal{R}$ consists of Clifford recovery operations (see Def. 5). If $\mathcal{W}$ teleports a logical qubit a distance $L>vT$ (13), then (i) $\mathcal{W}$ must fulfill the criteria of Lemma 8, (ii) the effective recovery channels (7) for the two logical basis operators (2.2) must be conditioned on different measurement outcomes, and (iii) the observables must satisfy
$\left[A_{1},\,A_{2}\right]=0\,,~{}~{}$ (41)
where $A_{1}$ and $A_{2}$ are the effective (involutory) operators (32) measured in the canonical-form protocol $\mathcal{W}$ (37).
The proof of Lemma 10 appears in App. H. That proof invokes many previous results, and particularly, the condition that $L>vT$. In addition to proving that $M=2$ is sufficient to realize $L>vT$ (13), the Lemma establishes a compatibility condition (41) on the canonical-form observables $A_{1,2}$ (32). Denoting by $A_{1}^{\prime}$ and $A_{2}^{\prime}$ the single-qubit observables measured in the naïve protocol $\mathcal{W}^{\prime}$ (which need not be in canonical form), we find
$\left[A_{1}^{\prime},\,U^{\dagger}_{12}\,A_{2}^{\prime}\,U_{12}\right]=0\,,~{}~{}$ (42)
if, in the Schrödinger picture, $\mathcal{W}^{\prime}$ involves the measurement of $A_{1}^{\prime}$, followed by the physical unitary channel $U_{12}$, followed by the measurement of $A_{2}^{\prime}$. Finally, we note that the two observables $A_{1,2}$ (either in the naïve protocol $\mathcal{W}^{\prime}$ or in the canonical-form protocol $\mathcal{W}$) must act on different sites, or they either fail to commute or fail to realize two distinct physical observables, so that Prop. 9 applies.
The compatibility condition (41) makes sense when one considers the information that can be extracted from the resource state $\lvert\Psi_{t}\rangle$ (36). Suppose that one measures $Z_{j}$ on site $j$. Measuring $Z_{j}$ again is gives the same outcome, while measuring $X_{j}$ or $Y_{j}$ gives a random outcome; in either case, no new information can be extracted. In the naïve protocol, the single-qubit measurements only extract information from the resource state if they act on distinct sites or if there are intervening multiqubit unitaries. For the canonical-form observables (32), one expects that at most $n$ independent outcomes can be extracted from $n$ qubits via measurements, provided that the $n$ observables mutually commute.
@@
Consider a standard teleportation protocol $\mathcal{W}$ (see Def. 6) in canonical form $\mathcal{W}=\mathcal{R}\,\mathcal{M}\,\mathcal{U}$ (37), where $\mathcal{U}$ is a local unitary (e.g., an FDQC) that obeys the bound (28) with Lieb-Robinson velocity $v$ and total duration (or depth) $T$, $\mathcal{M}$ involves the measurement of two compatible, involutory observables $A_{1}\neq A_{2}$, and $\mathcal{R}$ involves outcome-dependent operations that are useful in the sense of Lemma 10. Then the task distance $L=d(i,f)$ (13) obeys the bound
$L\leq v(3T-2)-2v\,\max_{\nu}\tau_{\nu}\,,~{}~{}$ (43)
where $\tau_{\nu}$ is the total duration of time evolution following the measurement of $A_{\nu}^{\prime}$ in the naïve protocol $\mathcal{W}^{\prime}$ (in the Schrödinger picture), and $T\geq 2$.
The proof of Lemma 11 appears in App. I, and relies on the fact that the light cones of both measured observables must intersect with that of the final-state logical operator $\sigma^{\nu}_{f}$ (at some site $\ell\approx i+2L/3$). The intuition is that measurements “reflect” (or “link up”) light cones Friedman et al. (2022a), leading to a threefold enhancement to $v$ in the $M=2$ bound (43) compared to $L\leq vT$ (28) for $M<2$.

### v3

@@ 4.1 Action of dilated channels
Here we consider the action of single-qubit measurements (2.3) and “Clifford” recovery operations (34) on the logical operators (4) in canonical-form protocols. We first derive an effective recovery channel for the logical operators $X^{\vphantom{*}}_{{\rm L}}$ and $Z^{\vphantom{*}}_{{\rm L}}$ in Lemma 7. We then consider the combination of measurements and feedback on the logical operators in Lemma 8. Both proofs extend to any $k\geq 1$.
We now present Lemma 7, which establishes an effective recovery channel $\mathcal{R}^{\,}_{\rm eff}$ for any $\mathcal{R}$ in a canonical-form standard teleportation protocol (see Defs. 4 and 6).
@@
$\mathcal{W}\cong\mathcal{W}^{\vphantom{\dagger}}_{\rm eff}=\mathcal{R}^{\vphantom{\dagger}}_{x}\,\mathcal{R}^{\vphantom{\dagger}}_{z}\,\mathcal{M}\,\mathcal{U}\,,~{}~{}$ (37)
where $\mathcal{R}^{\,}_{x}$ ($\mathcal{R}^{\,}_{z}$) act nontrivially only on $X^{\vphantom{*}}_{{\rm L}}$ ($Z^{\vphantom{*}}_{{\rm L}}$) as
$\mathcal{R}^{\vphantom{\dagger}}_{x}$ $=\mathds{1}\otimes\widetilde{P}^{(0)}_{r^{\,}_{x}}+Z^{\vphantom{*}}_{{\rm L}}\otimes\widetilde{P}^{(1)}_{r^{\,}_{x}}$ (38a)
$\mathcal{R}^{\vphantom{\dagger}}_{z}$ $=\mathds{1}\otimes\widetilde{P}^{(0)}_{r^{\,}_{z}}+X^{\vphantom{*}}_{{\rm L}}\otimes\widetilde{P}^{(1)}_{r^{\,}_{z}}\,,~{}~{}$ (38b)
where $r^{\,}_{\nu}\subset\widetilde{V}$ is a subset of Stinespring qubits determined by $\mathcal{R}$ (34), the projectors correspond to outcome parities of the cluster $r^{\,}_{\nu}$ (35), $\mathcal{R}^{\,}_{x}$ and $\mathcal{R}^{\,}_{z}$ effectively commute, and $\mathcal{R}^{\,}_{\nu}=\mathds{1}$ if $\mathcal{R}$ acts trivially on $\sigma^{\nu}_{\rm L}$.
The proof of Lemma 7 appears in App. E. The Lemma establishes (i) when and how a Clifford recovery operation $\mathcal{R}^{\,}_{s}$ (34) acts nontrivially on a logical operator $\Gamma$, (ii) that the effective recovery channel $\mathcal{R}^{\,}_{\rm eff}=\mathcal{R}^{\,}_{x}\,\mathcal{R}^{\,}_{z}$ (37) is equivalent to $\mathcal{R}$ in its action on all logicals $\Gamma$ (but not other operators), and (iii) that the recovery channels for different logical operators can be written in any order, and do not mix. For $k>1$ logical qubits, the pair of channels $\mathcal{R}^{\,}_{\nu}$ (37) are replaced by $2k$ channels $\mathcal{R}^{\,}_{n,\nu}$ for $n\in[1,k]$, in any order. We stress that both $\mathcal{M}$ and $\mathcal{U}$ in $\mathcal{W}$ (37) are unaltered compared to $\mathcal{W}=\mathcal{R}\,\mathcal{M}\,\mathcal{U}$ (31).
@@
Consider a standard teleportation protocol $\mathcal{W}$ in canonical form (see Defs. 4 and 6) involving the measurement of $M$ (single-qubit) observables $A^{\,}_{j}$ (32). Suppose that $\Gamma$ is a logical operator (e.g., $X^{\vphantom{*}}_{{\rm L}}$) acting on the final state $\lvert\Psi^{\,}_{T}\rangle$ (7). Then $\mathcal{W}$ fails to achieve teleportation (8) unless
$\left[A^{\vphantom{\prime}}_{j},\,\Gamma(t^{\vphantom{\prime}}_{j})\right]=0\,,~{}~{}$ (39)
@@
for all observables $A^{\,}_{j}$ (32) and all logical operators $\Gamma$ evolved to the Heisenberg time $t^{\,}_{j}$ immediately prior to the measurement of $A^{\,}_{j}$. Moreover, if $\mathcal{R}$ acts nontrivially on $\Gamma$, then the measurement channel acts as
$\Gamma(t^{\vphantom{\prime}}_{j-1})=\mathcal{M}^{\dagger}_{j}\,\Gamma(t^{\vphantom{\prime}}_{j})\,\mathcal{M}^{\vphantom{\dagger}}_{j}\cong\bar{A}^{\vphantom{\prime}}_{j}\,\Gamma(t^{\vphantom{\prime}}_{j})\,,~{}~{}$ (40)
@@
and trivially otherwise, where $\bar{A}^{\,}_{j}$ is the involutory part (33) of $A^{\,}_{j}$ (32) and $\cong$ becomes a genuine equality upon projecting onto the default initial Stinespring state $\lvert 0\rangle$.
@@
The proof of Lemma 8 appears in App. F. Simply put, Lemma 8 extends Lemma 7 to the combination of a measurements and recovery operations acting on involutory logical operators $\Gamma\in\boldsymbol{\Gamma}\subset\mathbf{P}^{\vphantom{*}}_{*}{\rm ph}$. For a given $\Gamma$, this combination either acts trivially, violates the assumption that $\Gamma(t)$ remains a logical operator (i.e., involutory), or attaches the involutory part $\bar{A}^{\,}_{j}$ (33) of the measured operator $A^{\,}_{j}$ to $\Gamma(t^{\,}_{j})$, where $t^{\,}_{j}$ is the (Heisenberg) time immediately prior to the $j$th measurement. Moreover, $A^{\,}_{j}$ must commute with $\Gamma(t^{\,}_{j})$. For $\mathcal{M}^{\,}_{j}$ to act nontrivially on $\Gamma$, an odd number of recovery Paulis $\chi^{\,}_{j,k}$ conditioned on the outcome $j$ must anticommute with the operators $\Gamma(\tau^{\,}_{j,k})$, where $\tau^{\,}_{j,k}$ is the (Heisenberg) time immediately prior to the channel $\mathcal{R}^{\,}_{j,k}$. This holds for each measured observable $A^{\,}_{j}$ and all logical operators $\Gamma$.
@@
Note that Lemma 8 holds for arbitrary (i.e., non-Clifford) $\mathcal{U}$ in the canonical-form protocol (31). The proof of Lemma 8 holds independently of any physical operations that act prior to the first measurement. Additionally, any physical unitary can be “pulled through” a measurement $\mathcal{M}^{\,}_{j}$ without affecting the result of Prop. 4, so that Lemma 8 continues to hold. Likewise, one can safely “pull through” any recovery operations $\mathcal{R}$ that act prior to some measurement (in the Schrödinger picture), to realize a canonical-form protocol (31).
@@
Summarizing the above, the action of any canonical-form standard teleportation protocol $\mathcal{W}$ on a logical operator $\Gamma\in\boldsymbol{\Gamma}$ is constrained by Lemmas 7 and 8. The former establishes that $\mathcal{R}$ is equivalent in its action on all $\Gamma\in\boldsymbol{\Gamma}$ to the same protocol $\mathcal{W}$ with $\mathcal{R}\to\mathcal{R}^{\,}_{\rm eff}=\prod_{n=1}^{k}\,\mathcal{R}^{\,}_{n,x}\,\mathcal{R}^{\,}_{n,z}$ (37), where $\chi^{\,}_{n,\nu}$ is the other logical operator for qubit $n$ or the identity. The latter establishes that each measured observable $A^{\,}_{j}$ must commute with every logical operator $\Gamma(t^{\,}_{j})$ at the Heisenberg time $t^{\,}_{j}$ immediately prior to $\mathcal{M}^{\,}_{j}$, and that $\mathcal{M}^{\,}_{j}$ attaches $\bar{A}^{\,}_{j}$ (33) to $\Gamma(t^{\,}_{j})$ if the effective channel $\mathcal{R}^{\,}_{\Gamma}$ (38) is nontrivial (i.e., $\chi^{\,}_{\Gamma}\neq\mathds{1}$).
@@ 4.2 At least two measurements are required
Consider a physical teleportation protocol $\mathcal{W}=\mathcal{R}\mathcal{M}\mathcal{U}$ (31) in canonical form (see Defs. 3 and 4), where $\mathcal{U}$ captures generic time evolution on $\mathcal{H}^{\,}_{\rm ph}$ (1) for duration $T$ and obeys the bound (28) with Lieb-Robinson velocity $v$; $\mathcal{M}$ realizes exactly one projective measurement of some (effectively) involutory observable $A$ (33); and $\mathcal{R}$ is an arbitrary recovery channel conditioned on the measurement’s outcome. If the protocol $\mathcal{W}$ teleports a single logical qubit a distance $L$ in time $T$, then $L\leq vT$.
The proof of Prop. 9 appears in App. G. Prop. 9 establishes that a single projective measurement cannot lead to a speedup compared to a protocol $\mathcal{W}^{\prime}$ with no measurements at all. Essentially, Prop. 1 is only fulfilled with $L>vT$ if the dilated channels attach distinct operators (via Lemma 8) to, e.g., the logical operators $X^{\vphantom{*}}_{{\rm L}}$ and $Z^{\vphantom{*}}_{{\rm L}}$; otherwise, $Y^{\vphantom{*}}_{{\rm L}}$ can only travel a distance $L\leq vT$.
@@
Consider a standard teleportation protocol $\mathcal{W}$ in canonical form $\mathcal{W}=\mathcal{R}\,\mathcal{M}\,\mathcal{U}$ (see Defs. 4 and 6), where $\mathcal{U}$ realizes unitary time evolution on $\mathcal{H}^{\,}_{\rm ph}$ for duration $T$ and obeys the bound (28), $\mathcal{M}$ involves two projective measurements, and $\mathcal{R}$ consists of Clifford recovery operations (see Def. 5). If $\mathcal{W}$ teleports a logical qubit a distance $L>vT$ (13), then (i) $\mathcal{W}$ must fulfill the criteria of Lemma 8, (ii) the effective recovery channels (38) for the two logical basis operators (4) must be conditioned on different measurement outcomes, and (iii) the observables must satisfy
$\left[A^{\,}_{1},\,A^{\,}_{2}\right]=0\,,~{}~{}$ (41)
where $A^{\,}_{1}$ and $A^{\,}_{2}$ are the effective (involutory) operators (32) measured in the canonical-form protocol $\mathcal{W}$ (37).
The proof of Lemma 10 appears in App. H. That proof invokes many previous results, and particularly, the condition that $L>vT$. In addition to proving that $M=2$ is sufficient to realize $L>vT$ (13), the Lemma establishes a compatibility condition (41) on the canonical-form observables $A^{\,}_{1,2}$ (32). Denoting by $A_{1}^{\prime}$ and $A_{2}^{\prime}$ the single-qubit observables measured in the naïve protocol $\mathcal{W}^{\prime}$ (which need not be in canonical form), we find
$\left[A_{1}^{\prime},\,U^{\dagger}_{12}\,A_{2}^{\prime}\,U^{\vphantom{\dagger}}_{12}\right]=0\,,~{}~{}$ (42)
if, in the Schrödinger picture, $\mathcal{W}^{\prime}$ involves the measurement of $A_{1}^{\prime}$, followed by the physical unitary channel $U^{\,}_{12}$, followed by the measurement of $A_{2}^{\prime}$. Finally, we note that the two observables $A^{\,}_{1,2}$ (either in the naïve protocol $\mathcal{W}^{\prime}$ or in the canonical-form protocol $\mathcal{W}$) must act on different sites, or they either fail to commute or fail to realize two distinct physical observables, so that Prop. 9 applies.
The compatibility condition (41) makes sense when one considers the information that can be extracted from the resource state $\lvert\Psi_{t}\rangle$ (36). Suppose that one measures $Z^{\vphantom{*}}_{j}$ on site $j$. Measuring $Z^{\vphantom{*}}_{j}$ again is gives the same outcome, while measuring $X^{\vphantom{*}}_{j}$ or $Y^{\vphantom{*}}_{j}$ gives a random outcome; in either case, no new information can be extracted. In the naïve protocol, the single-qubit measurements only extract information from the resource state if they act on distinct sites or if there are intervening multiqubit unitaries. For the canonical-form observables (32), one expects that at most $n$ independent outcomes can be extracted from $n$ qubits via measurements, provided that the $n$ observables mutually commute.
@@
Consider a standard teleportation protocol $\mathcal{W}$ (see Def. 6) in canonical form $\mathcal{W}=\mathcal{R}\,\mathcal{M}\,\mathcal{U}$ (37), where $\mathcal{U}$ is a local unitary (e.g., an FDQC) that obeys the bound (28) with Lieb-Robinson velocity $v$ and total duration (or depth) $T$, $\mathcal{M}$ involves the measurement of two compatible, involutory observables $A^{\,}_{1}\neq A^{\,}_{2}$, and $\mathcal{R}$ involves outcome-dependent operations that are useful in the sense of Lemma 10. Then the task distance $L=d(i,f)$ (13) obeys the bound
$L\leq v(3T-2)-2v\,\max_{\nu}\tau^{\,}_{\nu}\,,~{}~{}$ (43)
where $\tau^{\,}_{\nu}$ is the total duration of time evolution following the measurement of $A_{\nu}^{\prime}$ in the naïve protocol $\mathcal{W}^{\prime}$ (in the Schrödinger picture), and $T\geq 2$.
The proof of Lemma 11 appears in App. I, and relies on the fact that the light cones of both measured observables must intersect with that of the final-state logical operator $\sigma^{\nu}_{f}$ (at some site $\ell\approx i+2L/3$). The intuition is that measurements “reflect” (or “link up”) light cones [42], leading to a threefold enhancement to $v$ in the $M=2$ bound (43) compared to $L\leq vT$ (28) for $M<2$.